# Concise and Organized Perception Facilitates Reasoning in Large Language Models

### v4

@@ 4 Approach
Figure 3: Overview of the proposed COP with an example on DI-GSM (constructed from GSM8K [10]) with disordered and irrelevant premises. Green represents relevant premises [$p_{2},p_{3},p_{4},p_{6}$], black represents irrelevant premises [$p_{1},p_{5}$], and orange represents the question [$Q$]. Details of DI-GSM are listed in section5.1.
We present the Concise and Organized Perception (COP) reasoning approach, aiming to reconstruct concise and organized context as inputting to reduce the difficulty of model reasoning. Given a reasoning context containing plenty of premises $\mathcal{P}=\{p_{1},p_{2},...,p_{n}\}$, which may include relevant, irrelevant, or disorganized information, the task is to answer a question $\mathcal{Q}$ based on the given premises.
@@ 4.1 Capturing of locally-related premises
As mentioned previously, complex reasoning problems involve a set of premises $\mathcal{P}$. Imitating the process of human beings organizing thoughts, capturing of locally-related premises can be effectively achieved by employing directed edges to connect premises that bear relevance to one another. For example, an edge $p_{i}\rightarrow p_{j}$ connects premise $p_{i}$ to premise $p_{j}$ with the locally-related direction from $i$ to $j$. This approach facilitates a localized understanding of the relationship between the premises $p_{i}$ and $p_{j}$. For logical premises, especially modus ponens ($p_{a}$: if A then B; $p_{b}$: if B then C), capturing of locally-related premises can be performed by leveraging directed edges to connect each premise to premises whose consequents satisfy one or more of the conditions specified in the current premise. In the given case, premise $p_{b}$ can be directly connected to premise $p_{a}$. Besides logical premises, capturing of locally-related premises can be performed through semantic correlation, temporal correlation, etc. Figure 3(a) illustrates an instance consisting of premises with semantic correlation. Benefiting from LLMs’ powerful in-context learning and semantic understanding ability, our proposed method encourages LLMs to perform capturing locally-related premises by searching relevant premises for each premise in the given context, which can be prompted with few-shot examples. As shown in Figure 3(a), there are six premises in the original problems, which are disordered and contain two irrelevant premises. Through the understanding of semantic correlation, LLMs adeptly identified and correlated relevant premises within vast premises. For example, $p_{2}$ is connected to $p_{6}$ while $p_{1}$ is connected to "None" in Figure 3(a). After capturing locally-related premises for each premise, different pieces of locally-related premises can be integrated again through the connection directions between premises, ultimately forming a initial understanding of the original problem’s context. Figure 3(a) shows the four pieces of locally-related premises after integration, in which the relevance between premises is much clearer than that of the original input. Capturing of locally-related premises can be regarded as a preparatory step similar to the pre-processing that humans undertake before addressing complex tasks. The detailed prompts used in this step are listed in the Appendix A.3.
@@ 4.2 Generation of mind map
In section 4.1, locally-related premise fragments representing the preliminary structural understanding of the given reasoning context are captured. These captured pieces are independent and cannot be directly merged. Serving as an anchor point, the query question functions as a connecting bridge. Upon receiving a question, relevant clues can be identified among the pieces of locally-related premises, allowing for the creation of a tree-like mind map structure relevant to the question. This process effectively eliminates irrelevant information and presents a coherent global understanding of the relevant information in an ordered manner.
Specifically, COP encourages LLMs to find all premises centered around the question $\mathcal{Q}$, which can also be prompted with few-shot examples. The detailed prompts used in this step are listed in the Appendix A.3. As shown in Figure 3(b), two relevant pieces of locally-related premises are involved, and another two irrelevant ones are discarded. Although "If Sara already had \$10 saved before Sara started babysitting" is a logical description of the question, COP can still correlate it to other pieces of locally-related premises, ultimately forming a holistic and ordered structure. Once we find the relevant pieces of locally-related premises, we perform a $D$-depth searching starting from each of them to avoid reasoning loops, where $D$ is the max reasoning depth. In this way, a tree-like mind map based on the given question is constructed. Generating a tree-like mind map structure based on questions as anchor points is a crucial step in our method that facilitates the development of a structured arrangement centered around the core question while eliminating irrelevant information. Such a strategy provides a foundational guarantee for subsequent concise and organized context reconstruction, aligning with human problem-solving preferences.
@@ 4.3 Context reconstruction
5 Experiments
5.1 Datasets
(1) ProofWriter [35] is a commonly used logical reasoning dataset and contains five subsets, named $d5$, $d3$, $d2$, $d1$ and $d0$ respectively. $dx$ part requires $\leq x$ hops for reasoning. We randomly sampled 600 examples in each part. (2) PrOntoQA is a synthetic logical reasoning dataset [^2] https://github.com/asaparov/prontoqa/tree/v1 and we use three parts $hop5$, $hop3$, and $hop1$ for testing. $hopx$ part requires $x$ hops for reasoning. We randomly sampled 500 examples in each part. (3) PrOntoQA-OOD is another synthetic logical reasoning dataset containing different types of premises. We used the generated data file generated_ood_data.zip based on the open-source repository [^3] https://github.com/asaparov/prontoqa, and randomly selected 300 samples from the original $hop2$ part for testing. (4) FOLIO [18] is a real-world dataset for logical reasoning, written in highly natural wordings. We randomly sampled 100 examples for testing. (5) DI-GSM is a constructed dataset containing dirordered and irrelevant information. Referring to [7], we first select GSM8K [10] test problems with at least five sentences in the problem statements and shuffle the sentences. Besides, we randomly add 2 to 3 irrelevant statements to the questions. The final testing data in DI-GSM contains 132 problems.
5.2 Experimental results
5.2.1 Performance comparison with state-of-the-art methods
In this section, we perform a thorough comparison between our proposed method and the existing state-of-the-art methods (Standard Few-Shot, CoT [41], SI [11], LOGIC-LM [30] and LAMBADA [25]). As typical methods of teaching model how to plan, SI [11] suggests alternating between selection and inference to generate reasoning steps based on forward chaining, while LAMBADA [25] employs a more explicit manner to introduce backward chaining for high-level proof planning. Since our approach naturally comes from a different perspective easy to plan, it can be seamlessly combined with other popular methods, such as SI or LAMBADA, to further enhance their performance. Unless otherwise specified, all the experimental results of COP are based on GPT-3.5-Turbo [29].
Table 1: Comparison of label accuracy on ProofWriter, PrOntoQA and PrOntoQA-OOD.
Datasets/ ProofWriter PrOntoQA
@@ 5.2.1 Performance comparison with state-of-the-art methods
Methods d5 d3 d2 d1 d0 average 5-hop 3-hop 1-hop average -
Standard 41.67 49.83 51.00 55.50 63.67 52.33 49.60 52.00 65.60 55.73 43.33
CoT 53.50 61.17 61.33 62.33 62.83 60.23 69.80 74.20 86.20 76.73 85.67
SI 46.00 51.00 56.00 61.00 97.00 62.20 45.00 52.00 97.00 64.67 -
LogicLM 70.11 - - - - - 93.20 - - - -
LAMBABDA 72.00 82.00 87.00 90.00 98.00 85.80 96.00 99.00 98.00 97.67 38.33
COP 88.67 90.67 91.43 92.50 98.50 91.72 99.20 99.60 100.00 99.60 94.00
@@
Table 1 shows the results on ProofWriter, ProntoQA, and PrOntoQA-OOD. ProofWriter and ProntoQA contain various reasoning depths. The results of SI and LAMBADA are taken from [25]. COP consistently achieves the highest label accuracy across all experimental settings. Notably, COP outperforms SOTA methods by a large margin on the hardest Depth-5 subset of ProofWriter. It shows a remarkable 65.74% relative improvement compared to CoT and 23.15% compared to LAMBADA, which demonstrates the effectiveness of COP. Table 2 shows the results on FOLIO and DI-GSM. COP outperforms CoT and LogicLM, while LogicLM is not able to work on DI-GSM, and LAMBADA is not able to work on these two datasets, demonstrating the efficacy of COP. The results of LogicLM on FOLIO are reproduced, and the logic program errors in LogicLM are considered as the failure cases in our report. Besides, COP improves the performance while reducing the cost, and the detailed comparison of average token usage is listed in  A.2.3.
Table 2: Test on FOLIO and DI-GSM.
Methods FOLIO DI-GSM
Standard 23.00 46.97
CoT 38.00 44.70
LogicLM 34.00 -
COP 47.00 53.79
In the next sections, we will analyze COP from several aspects, including concise and organized perception analysis from the perspective of information flow ( 5.2.2), error analysis caused by COP itself ( 5.2.3), error analysis when combined with other methods ( 5.2.4). In addition, COP consistently achieves high accuracy using different LLMs (including Qwen2-72B-Instruct, Llama-3-70B/8B-Instruct, and Llama-2-13B-Chat), and the detailed results are listed in  A.2.1.
5.2.2 Concise and organized perception analysis
As analyzed in section3, disordered and irrelevant content has a negative impact on reasoning from the perspective of information flow. The proposed COP seeks to enhance reasoning performance and simplify the reasoning process by systematically removing irrelevant information and reorganizing the inputs. Thus, we analyze the performance of COP from two aspects: the degree of removal of irrelevant information and the order of input. Firstly, each sample in the DI-GSM dataset contains 2 to 3 irrelevant statements to the questions, for a total of 322 irrelevant statements. After applying COP, only nine statements have not been removed, demonstrating the success of concise perception in COP. Secondly, the statements of questions are randomly shuffled in DI-GSM. We use Kendall tau distance to measure the order difference of statements between the shuffled data and the original data. Kendall tau distance ranges from -1 to 1. The larger the Kendall tau distance, the more relevant it is. Before applying COP, the average Kendall tau distance is -0.0465, indicating no strong correlation between the statement order in the shuffled data and the original data. After applying COP, the average Kendall tau distance increases to 0.4268, demonstrating the success of organized perception in COP.
@@ 5.2.3 Failure case analysis
In this section, we focus on the errors caused by COP itself to study its possible flaws. We compare the results of COP on DI-GSM and CoT on DI-GSM with original concise and organized inputs based on GPT-3.5-turbo. The accuracy of COP is 53.79 (71/132), and the accuracy of CoT is 68.94 (91/132). We analyze the differences in the results and find that COP gets 15 more questions correct compared to CoT, and COP fails 35 more questions than CoT. After applying COP, the order of statements in the 15 correct questions is different from the original order, and the changed order generated by COP is more suitable for reasoning, demonstrating the success of organized perception in COP. The error types of the 35 failure cases are shown in Table 3. There are 10 cases where the order of the premises caused failure in reasoning. There are 13 cases where the premises are not connected to other premises in the given context. There are 2 cases where the premises are not output in the step of capturing of locally-related premises. These failure cases are attributed to the fact that capturing of locally-related premises is driven by LLMs, and it is difficult for the method to ensure that the generated connections between premises are completely accurate. Similarly, for the remaining 10 cases, some key relevant premises were discarded when generating mind maps, causing failure in reasoning. We leave the failure caused by these error types as future work.
@@
Step Capturing of Locally-related Premises Generation of Mind Map
Type Connection Order Connection Output Connection None Connection Discard
@@
5.2.4 Is COP beneficial to other methods ?
Based on easy to plan, our COP can be seamlessly combined with methods that teach models how to plan, such as LAMBADA. Since LAMBADA cannot work on two real-world datasets, FOLIO and DI-GSM, we use ProofWriter for testing. The performance of LAMBADA and LAMBADA+COP on the ProofWriter $d_{5}$ subset with different inference depths are listed in Figure 5 (a). All the results are based on the LAMBADA code we reproduced, and the base model of this experiment is GPT-3.5-turbo. Compared with the original LAMBADA method, the performance of LAMBADA+COP under different inference depths is improved, proving the effectiveness of COP. In addition, Figure 5(b)(c) shows the proportion of correct reasoning and the proportion of different types of incorrect reasoning. We randomly selected 100 test samples from the ProofWriter $d_{5}$ subset to manually check the error types of incorrect reasoning examples. As shown in the figure, equipped with COP, the proportion of selection errors (including fact check and rule selection modules in LAMBADA) drops significantly. The proportion of goal decomposition errors and sign agreement errors (goal decomposition and sign agreement modules are unaffected by context redundancy and disorder) are almost unchanged, proving that our COP can improve the success rate of other methods in the steps that are affected by context redundancy and disorder.
Figure 5: (a) Comparison of LAMBADA and LAMBADA+COP on Proofwriter. (b) The proportions of different error reasons of LAMBADA. (c) The proportions of different error reasons of LAMBADA + COP. There are four steps in LAMBADA, which are Fact Check step, Rule Selection step, Goal Decomposition step and Sign Agreement step. "Incorrect-Selection" means that LAMBADA fails in Fact Check or Rule Selection steps. "Incorrect-Goal" means that LAMBADA fails in Goal Decomposition steps. "Incorrect-Sign" means that LAMBADA fails in Sign Agreement steps.
6 Conclusion and future work
@@
We believe our crucial insight on the proposal of easy-to-plan method has broader implications. However, for more general reasoning tasks, performing robust capturing of locally-related premises and generating more appropriate tree-like mind map structures require further exploration. We plan to address these in future research.

### v5

@@ 4 Approach
Figure 4: Overview of the proposed COP with an example on DI-GSM (constructed from GSM8K (Cobbe et al., 2021)) with disordered and irrelevant premises. Green represents relevant premises [$p_{2},p_{3},p_{4},p_{6}$], black represents irrelevant premises [$p_{1},p_{5}$], and orange represents the question [$Q$]. Details of DI-GSM are listed in section5.1.
We present the Concise and Organized Perception (COP) reasoning approach, aiming to reconstruct concise and organized context as inputting to reduce the difficulty of model reasoning. Given a reasoning context containing plenty of premises $\mathcal{P}=\{p_{1},...,p_{n}\}$, which may include relevant, irrelevant, or disorganized information, the task is to answer a question $\mathcal{Q}$ based on the given premises.
@@ 4.1 Capturing of locally-related premises
As mentioned previously, complex reasoning problems involve a set of premises $\mathcal{P}$. Imitating the process of human beings organizing thoughts, capturing of locally-related premises can be effectively achieved by employing directed edges to connect premises that bear relevance to one another. For example, an edge $p_{i}\rightarrow p_{j}$ connects premise $p_{i}$ to premise $p_{j}$ with the locally-related direction from $i$ to $j$. This approach facilitates a localized understanding of the relationship between the premises $p_{i}$ and $p_{j}$. For logical premises, especially modus ponens ($p_{a}$: if A then B; $p_{b}$: if B then C), capturing of locally-related premises can be performed by leveraging directed edges to connect each premise to premises whose consequents satisfy one or more of the conditions specified in the current premise. In the given case, premise $p_{b}$ can be directly connected to premise $p_{a}$. Besides logical premises, capturing of locally-related premises can be performed through semantic correlation, temporal correlation, etc. Figure 4(a) illustrates an instance consisting of premises with semantic correlation. Benefiting from LLMs’ powerful in-context learning and semantic understanding ability, our proposed method encourages LLMs to perform capturing locally-related premises by searching relevant premises for each premise in the given context, which can be prompted with few-shot examples. As shown in Figure 4(a), there are six premises in the original problems, which are disordered and contain two irrelevant premises. Through the understanding of semantic correlation, LLMs adeptly identified and correlated relevant premises within vast premises. For example, $p_{2}$ is connected to $p_{6}$ while $p_{1}$ is connected to "None" in Figure 4(a). After capturing locally-related premises for each premise, different pieces of locally-related premises can be integrated again through the connection directions between premises, ultimately forming a initial understanding of the original problem’s context. Figure 4(a) shows the four pieces of locally-related premises after integration, in which the relevance between premises is much clearer than that of the original input. The detailed prompts used in this step are listed in the Appendix A.7.
@@ 4.2 Generation of mind map
The above captured locally-related premise fragments are independent and cannot be directly merged. Serving as an anchor point, the query question functions as a connecting bridge. Upon receiving a question, relevant clues can be identified among the pieces of locally-related premises, allowing for the creation of a tree-like mind map structure relevant to the question. This process effectively eliminates irrelevant information and presents a coherent global understanding of the relevant information in an ordered manner.
Specifically, COP encourages LLMs to find all premises centered around the question $\mathcal{Q}$, which can also be prompted with few-shot examples. The detailed prompts used in this step are listed in the Appendix A.7. As shown in Figure 4(b), two relevant pieces of locally-related premises are involved, and another two irrelevant ones are discarded. Although "If Sara already had \$10 saved before Sara started babysitting" is a logical description of the question, COP can still correlate it to other pieces of locally-related premises, ultimately forming a holistic and ordered structure. Once we find the relevant pieces of locally-related premises, we perform a $D$-depth searching starting from each of them to avoid reasoning loops, where $D$ is the max reasoning depth. In this way, a tree-like mind map based on the given question is constructed. Generating a tree-like mind map structure based on questions as anchor points is a crucial step in our method that facilitates the development of a structured arrangement centered around the core question while eliminating irrelevant information. Such a strategy provides a foundational guarantee for subsequent concise and organized context reconstruction.
@@ 4.3 Context reconstruction
Table 1: Comparison of accuracy with various LLMs. "X" means the method has difficulty in running on DI-GSM.
Models GPT-4o Claude-3-5-Sonnet Llama-3-70B-Instruct GPT-3.5-Turbo
Datasets/Methods FOLIO DI-GSM
ProofWriter
-d5
FOLIO DI-GSM
ProofWriter
-d5
DI-GSM
ProofWriter
-d5
ProofWriter
-d5
PrOntoQA
-5hop
@@
Standard 58.00 80.30 61.33 64.00 79.55 65.00 76.52 55.00 41.67 49.60 43.33
CoT 60.00 83.33 77.33 65.00 85.61 65.67 72.73 67.50 53.50 69.80 85.67
IRRE 62.00 83.33 76.50 66.00 84.09 66.67 77.27 66.17 52.17 77.20 80.67
S2A 63.00 74.24 73.33 67.00 80.30 57.33 73.48 61.00 43.00 56.00 74.67
SI - X - - X - X - 46.00 45.00 -
LogicLM 56.00 X 68.33 55.00 X 77.83 X 70.17 51.40 51.58 -
LAMBABDA - X - - X - X - 72.00 96.00 38.33
COP 65.00 84.85 96.50 68.00 86.36 94.50 78.79 83.50 88.67 99.20 94.00
Table 2: Comparison of accuracy based on Llama-3-8B-Instruct and Qwen2-72B-Chat.
Models Llama-3-8B-Instruct Qwen2-72B-Chat
Datasets/Methods DI-GSM ProofWriter-d5 DI-GSM ProofWriter-d5
Standard 35.61 44.17 71.21 46.67
CoT 43.18 44.33 78.79 63.83
IRRE 50.00 43.17 71.97 63.17
S2A 40.91 45.00 75.76 56.33
LogicLM X 31.83 X 48.25
COP 52.27 77.67 80.30 75.67
Table 3: Comparison of average inference calls and token numbers on FOLIO and ProofWriter.
FOLIO
Method Calls Prompt-tokens Total-tokens
LogicLM 4 6204.3957 7281.5731
COP 3 3801.9971 4104.9285
ProofWriter-5Hop
Method Calls Prompt-tokens Total-tokens
LAMBADA 24.71 19200.05 21922.77
COP 3.87 2004.97 2440.26
5 Experiments
5.1 Datasets and compared methods
For Real-world Settings, we perform experiments on FOLIO and DI-GSM. For Synthetic Settings, we perform experiments on ProofWriter, PrOntoQA and PrOntoQA-OOD. Details about datasets can be found in  A.1. We perform a thorough comparison between our proposed method and the existing state-of-the-art methods (Standard Few-Shot, CoT (Wei et al., 2022), IRRE (Shi et al., 2023), S2A (Weston and Sukhbaatar, 2023), SI (Creswell et al., 2023), LogicLM (Pan et al., 2023) and LAMBADA (Kazemi et al., 2023)). Details of the compared methods can be found in  A.2.
5.2 Experimental results
5.2.1 Results Analysis
@@
Comparison with SOTA. Table 1 and Table 2 show the results on different datasets. For SI and LAMBADA, as they just report results based on GPT-3.5-Turbo(Kazemi et al., 2023), we also conduct experiments on GPT-3.5-Turbo for fair comparison. COP consistently achieves the highest accuracy across all experimental settings. Notably, COP outperforms SOTA methods by a large margin on the hardest Depth-5 subset of ProofWriter. It shows a remarkable 65.74% relative improvement compared to CoT and 23.15% compared to LAMBADA, which demonstrates the effectiveness of COP. LogicLM converts natural language questions into first-order logical symbolics and performs symbolic reasoning to accomplish logical reasoning tasks. All problems in DI-GSM are mathematical problems, making it difficult for LogicLM to represent them using first-order logic. As shown Table 1 and Table 2, COP outperforms CoT and LogicLM, while LogicLM is not able to work well on real-world setting DI-GSM (marked as "X"), demonstrating the efficacy of COP. Details of mark "X" are listed in  A.2. The results of LogicLM on FOLIO are reproduced, and the logic program errors in LogicLM are considered as the failure cases in our report. ProofWriter and ProntoQA contain various reasoning depths, detailed results on different reasoning depths are listed in  A.5.1.
Generalizability. COP consistently achieves high accuracy using different LLMs including GPT-4o, Claude3.5-Sonnet, Llama-3-70B/8B-Instruct, Qwen2-72B-Instruct, crossing both synthetic setting and real-world setting (FOLIO and DI-GSM). Additional results on more different size open-source and close-source LLMs including Qwen1.5-72B-Chat, Llama-2-13B-Chat, mistral-7b-instruct-v0.3, Gemini-1.0-Pro and GPT-4o-mini can be found in  A.5.2.
Efficiency. Besides, as shown in Table 3, COP improves the performance while reducing the cost. COP utilize two efficient strategies (concise and organized) to identify ordered and related information, reducing the token consumption of irrelevant information and longer reasoning path caused by disorder during reasoning. The detailed comparison of average token usage can be found in  A.5.3.
Ablation study. We conduct experiments on DI-GSM to analyze the effectiveness of the components in COP. Results are in Table 4. On the one hand, without the mind map generation step in COP, the accuracy drops (only 43.94), proving the importance of identifying relevant premises around the given question, which is a well-designed step in our COP. One the other hand, we compare with two approaches (Zhao et al., ; Guo et al., 2023) instead of our mind map generation step. COP is also the best and effective. Details can be found in  A.2.
Table 4: Analysis of the components in COP.
COP
w/o mind map
generation
w/ m1
 (Zhao et al., )
w/ m2
 (Guo et al., 2023)
53.79 43.94 41.67 43.18
@@ 5.2.2 Failure case analysis
COP. In this section, we focus on the errors caused by COP itself to study its possible flaws. We compare the results of COP on DI-GSM and CoT on DI-GSM with original concise and organized inputs based on GPT-3.5-turbo. The accuracy of COP is 53.79 (71/132), and the accuracy of CoT is 68.94 (91/132). We analyze the differences in the results and find that COP gets 15 more questions correct compared to CoT, and COP fails 35 more questions than CoT. After applying COP, the order of statements in the 15 correct questions is different from the original order, and the changed order generated by COP is more suitable for reasoning, demonstrating the success of organized perception in COP. The error types of the 35 failure cases are shown in Table 5. There are 10 cases where the order of the premises caused failure in reasoning. There are 13 cases where the premises are not connected to other premises in the given context. There are 2 cases where the premises are not output in the step of capturing of locally-related premises. These failure cases are attributed to the fact that capturing of locally-related premises is driven by LLMs, and it is difficult for the method to ensure that the generated connections between premises are completely accurate. Similarly, for the remaining 10 cases, some key relevant premises were discarded when generating mind maps, causing failure in reasoning. We present several case studies in  A.6 for clearer understanding. We leave the failure caused by these error types as future work. Besides, our approach naturally comes from a different perspective easy to plan, it can be seamlessly combined with other popular methods, such as LAMBADA, to further enhance their performance.
@@
Step
Capturing of Locally
-related Premises
Generation of
Mind Map
Type
Connection
Order
Connection
Output
Connection
None
Connection
Discard
@@
Figure 6: (a) Comparison of LAMBADA and LAMBADA+COP on Proofwriter. (b) The proportions of different error reasons of LAMBADA and LAMBADA+COP. There are four steps in LAMBADA, which are Fact Check, Rule Selection, Goal Decomposition and Sign Agreement. "Incorrect-Selection/Goal/ Sign" means that LAMBADA fails in Fact Check or Rule Selection/Goal Decomposition/Sign Agreement steps.
COP combined with LAMBADA. Based on easy to plan, our COP can be seamlessly combined with LAMBADA. The performance of LAMBADA and LAMBADA+COP on ProofWriter $d_{5}$ subset are listed in Figure 6 (a). Compared with the original LAMBADA method, the performance of LAMBADA+COP under different inference depths is improved, proving the effectiveness of COP. In addition, Figure 6(b) proves that COP can improve the success rate of LAMBADA in the steps that are affected by context redundancy and disorder. Details can be found in  A.5.6.
6 Conclusion
@@
Limitations
In this work, we focus on two issues including distractibility and disorder. And we focus on the task characterized by plenty of premises, which may contain disordered and irrelevant content. Tasks in which premises do not have an obvious order and do not involve irrelevant information are outside the scope of this paper.
We believe our crucial insight on the proposal of easy-to-plan method has broader implications. However, as in error analysis, COP may fails on some examples, but the overall performance improves across different models (results are listed in Main text and Appendix). For more general reasoning tasks, performing robust capturing of locally-related premises and generating more appropriate tree-like mind map structures require further exploration. We plan to address these issues in future research.